The width of a rectangle is 3 meters more than one fourth its length. The perimeter is 10 meters more than twice its length. Find the length and width.
Answer by stanbon(75887)

The width of a rectangle is 3 meters more than one fourth its length. The perimeter is 10 meters more than twice its length. Find the length and width.
---
Let length be "x":
Then width = "(x/4)+3"
----
Perimeter = 2(length + width)
10 = 2(x + (x/4) + 3))
5 = (5/4)x + 3
(5/4)x = 2
x = 2(4/5)
x = 8/5 meters (length)
---------------
(x/4)+3 = (2/5)+3 = 3 2/5 meters (width)
============================================
Cheers,
Stan H.